A bond has a face value of ​$1200 and a 10​% coupon​ rate, its current price is ​$1140​ and it is expected to increase to ​$1150 next year.
Answer #1

Calculating the current yield

Coupon = 10% of $ 1,200 = $ 120

Current price = $ 1,140

Now calculating the capital gain.

Capital gain = P1 - P0

Capital gain = $ 1,150 - $ 1,140 = $ 10

Expected rate of return = current yield + capital gain

Expected Rate of Return = 10.53 + 0.88 = 11.41 %

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me, please help. Thank you.